mbkpizzalover:

my recent marketing idea, the miracle algae growing kit, has been remarkably successful, with monthly sales growing by 4% every 6 months over the past 4 years. Assuming that i sold 100 kits the first month, how many kits did I sell in the first month of the year?

mbkpizzalover:

Pls help

jhonyy9:

like a first step using the rule of 3 simple you can calcule the sales growing for one month 4% for 6 months x% for 1 month x = (1*4%)/6 = (4/100)/6 = ?